OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of –3?

OpenStudy (anonymous):

Hey, I just wanted to know that you are in the Physics Section and that your question deals with Mathematical Concepts. It was most likely a mistake, but here's the link to the Code-Of-Conduct, you should probably read it to remember to post your question in the right section: http://openstudy.com/code-of-conduct And, so that you can post this question in the correct section, here is the link to the Mathematics Section: http://openstudy.com/study#/groups/Mathematics I also hope this doesn't come off as rude, because it's meant to be informative :)

OpenStudy (shamim):

Ok. The equation of a straight line is y=mx+c -2=-3*1+c c=?

OpenStudy (shamim):

Put these value of c and m in my first equation. U will get answer